# Composed Image Retrieval for Remote Sensing

### v2

@@ 4.2 Experimental results
Comparison with baselines As shown in Table 2, WeiCom outperfoms both unimodal (“Text”, “Image”) and multimodal (“Text & Image”) baselines by a large margin. In particular, it outperforms the second best by 8.50% mAP using CLIP and 11.66% mAP using RemoteCLIP on average.
@@
Text 14.14 4.83 3.58 4.38 6.30 6.22 6.58
Image 11.80 8.32 13.49 13.50 6.30 15.76 11.53
Text & Image 19.59 11.02 15.87 13.77 7.82 21.38 14.91
$\textsc{WeiCom}_{\lambda=0.5}$ 41.15 17.45 16.49 9.24 18.00 23.97 21.05
$\textsc{WeiCom}_{\lambda=0.3}$ 40.71 20.97 22.07 12.07 18.40 26.22 23.41
@@
Text 11.89 8.87 22.16 12.49 12.56 24.12 16.99
Image 11.72 6.62 15.11 9.29 5.41 15.18 11.19
Text & Image 19.84 10.01 18.45 10.56 6.23 19.63 14.85
$\textsc{WeiCom}_{\lambda=0.5}$ 40.08 31.45 39.94 14.27 14.14 29.78 28.28
$\textsc{WeiCom}_{\lambda=0.6}$ 38.20 31.59 41.56 14.79 14.53 31.24 28.65

### v3

@@ 4.2 Experimental results
Comparison with baselines As shown in Table 2, WeiCom outperfoms both unimodal (“Text”, “Image”) and multimodal (“Text & Image”) baselines by a large margin. In particular, it outperforms the second best by 8.95% mAP using CLIP and 15.14% mAP using RemoteCLIP on average.
@@
Text 13.47 4.83 3.58 4.38 3.31 6.22 5.97
Image 14.66 8.32 13.49 13.50 7.84 15.76 12.26
Text & Image 23.13 11.02 15.87 13.77 10.13 21.38 15.88
$\textsc{WeiCom}_{\lambda=0.5}$ 46.08 17.45 16.49 9.24 18.15 23.97 21.90
$\textsc{WeiCom}_{\lambda=0.3}$ 46.74 20.97 22.07 12.07 20.96 26.22 24.83
@@
Text 10.75 8.87 22.16 12.49 8.25 24.12 14.44
Image 14.40 6.62 15.11 9.29 6.99 15.18 11.27
Text & Image 23.67 10.01 18.45 10.56 7.97 19.63 15.05
$\textsc{WeiCom}_{\lambda=0.5}$ 43.68 31.45 39.94 14.27 20.51 29.78 29.94
$\textsc{WeiCom}_{\lambda=0.6}$ 41.04 31.59 41.56 14.79 20.79 31.24 30.19